The electron gun in a television tube is used to accelerate electrons with mass 9.109
Answer #1

a. apply Grav force = elec force

ma = Eq

here use V^2 = 2aS

so a = v^2/2S

a = (1e7*1.e7)/(2* 0.014)

a = 3.57 *10^15 m/s^2

for electron

Eelectric field E = 9.11e-31 * 3.57 e15/(1.6 e-19)

E = 2.032 *10^4 N/C
Answer #2

vo = 0

v = 1x 10^7 m/s

d = 1.4 cm = 1.4 x 10^-2 m

v^2 = vo^2 + 2ad

(1 x 10^7)^2 = 0 + 2 a (1.4 x 10^-2)

10^14 = 2.8 x 10^-2 a

a= 3.5714*10^15

F =ma

F =9.109*10^-31 *3.5714*10^15

F= 3.2532*10^-15 N

E= F/q

E =3.2532*10^-15/1.6*1.6-19

E= 20332.589 N/C
Answer #3

The electron gun in a television tube is used to accelerate electrons with mass 9.109
Answer #4

I think that 32N/C
Answer #5

My answer 43N/C